# thanks: These authors equally contributed to this workthanks: These authors equally contributed to this workthanks: A professor emeritus

### v1

@@ Abstract
Discovering and categorizing quantum orders in mixed many-body systems are currently one of the most important problems. Target model in this study is an extended version of the cluster model in one dimension with $Z_{2}\otimes Z_{2}$ symmetry, and we investigate effects of decoherence applied to the ground state of the model, focusing on the symmetry aspect. By using a scheme that we propose, a strong symmetry protected topological (SPT) mixed state and double average SPT (ASPT) state are constructed through the pure gapless SPT order and the domain-wall duality. Among them, the double ASPT is categorized by coexisting orders, i.e., a strong-to-weak spontaneous symmetry breaking and ASPT defined by the remaining weak and strong symmetries. We make use of the doubled Hilbert space formalism for the construction scheme. We numerically demonstrate the emergence of the two mixed SPT states and find that a phase transition takes place between them tuned by the strength of decoherence. Suitable multiple-decoherence channel applied to SPT states gives a broad possibility to induce rich ASPTs, possessing non-trivial internal entanglement properties from the view of doubled Hilbert space formalism.
@@ I Introduction
The rest of this paper is organized as follows. In Sec. II, we review the construction scheme and protocol of the gSPT. In Sec. III, based on this scheme, we propose its extensive application to a mixed state, where our target system under decoherence is introduced. We explain the doubled Hilbert space formalism and show the interpretation of the decoherence in this formalism. We expect the emergence of non-trivial mixed SPTs. In Sec. IV, we show the systematic numerical demonstrations by using the MPS and the filtering to the MPS for various decoherence parameters. Here, we verify the properties of the non-trivial mixed SPTs and a phase transition between them. Section V is devoted for summary and conclusion.
@@ III Construction of non-trivial mixed SPT: Double ASPT
    Possibility of gapless ASPT mixed state: In the pure state system for $J_{zz}=1$, gSPT, $|{\rm gSPT}\rangle$, emerges. We predict its mixed state counter part since in the $\sigma$-spin system before DW transformation, the ^$\hat{\mathcal{E}}_{\sigma^{x}}(p_{x})$ and ^$\hat{\mathcal{E}}_{\sigma^{z}\sigma^{z}}$ induce the SWSSB phase transition as verified in [38]. We expect that the target mixed state $|\rho^{DW}_{D}\rangle\rangle$ obtained by the DW duality transformation inherits its phase diagram. At the critical decoherence strength $p_{zz}=p^{c}_{zz}$ (if it exists), the decohered state $|\rho^{DW}_{D}\rangle\rangle$ can be regarded as a mixed-state counterpart to gSPT [39, 40], namely gapless ASPT.
@@ IV.3 Physical observables
The first ones are $\sigma$-spin correlators to identify the $Z^{\sigma}_{2}$ SWSSB state. The one of them is the reduced susceptibility of Rényi-2 correlator,
@@
Another observable is the reduced susceptibility of the $\sigma^{z}\sigma^{z}$-correlator of the upper chain,
@@ IV.4 Numerical results
V Summary and conclusion

### v2

@@ Abstract
Discovering and categorizing quantum orders in mixed many-body systems are currently one of the most important problems. Target model in this study is an extended version of the cluster model in one dimension with $Z_{2}\otimes Z_{2}$ symmetry, and we investigate effects of decoherence applied to the ground state of the model, focusing on the symmetry aspect. By using a scheme that we propose, a strong symmetry protected topological (SPT) mixed state and double average SPT (ASPT) state are constructed through the pure gapless SPT order and the domain-wall duality. Among them, the double ASPT is categorized by coexisting orders, i.e., a strong-to-weak spontaneous symmetry breaking and ASPT defined by the remaining weak and strong symmetries. We make use of the doubled Hilbert space formalism for the construction scheme. We numerically demonstrate the emergence of the two mixed SPT states and find that a phase transition occurs between them tuned by the strength of decoherence. Finally, we discuss the coexistence of SPT and SWSSB in the double SPT state from the view point of symmetrically invertible property, and comment on the classification of ASPT proposed recently. Suitable multiple-decoherence channel applied to SPT states gives a broad possibility to induce rich ASPTs, possessing non-trivial internal entanglement properties from the view of doubled Hilbert space formalism.
@@ I Introduction
The rest of this paper is organized as follows. In Sec. II, we review the construction scheme and protocol of the gSPT. In Sec. III, based on this scheme, we propose its extensive application to a mixed state, where our target system under decoherence is introduced. We explain the doubled Hilbert space formalism and show the interpretation of the decoherence in this formalism. We expect the emergence of non-trivial mixed SPTs. In Sec. IV, we show the systematic numerical demonstrations by using the MPS and the filtering to the MPS for various decoherence parameters. Here, we verify the properties of the non-trivial mixed SPTs and a phase transition between them. In Sec. V, we briefly discuss the coexistence of SPT and SWSSB in the double SPT state from the view point of symmetrically invertible property, and comment on the classification of ASPT proposed recently. Section VI is devoted for summary and conclusion.
@@ III Construction of non-trivial mixed SPT: Double ASPT
    Possibility of gapless ASPT mixed state: In the pure state system for $J_{zz}=1$, gSPT, $|{\rm gSPT}\rangle$, emerges. We predict its mixed state counter part since in the $\sigma$-spin system before DW transformation, the ^$\hat{\mathcal{E}}_{\sigma^{x}}(p_{x})$ and ^$\hat{\mathcal{E}}_{\sigma^{z}\sigma^{z}}$ induce the SWSSB phase transition as verified in [38]. We expect that the target mixed state $|\rho^{DW}_{D}\rangle\rangle$ obtained by the DW duality transformation inherits its phase diagram. At the critical decoherence strength $p_{zz}=p^{c}_{zz}$ (if it exists), the decohered state $|\rho^{DW}_{D}\rangle\rangle$ can be regarded as a mixed-state counterpart to gSPT [39, 40], namely gapless ASPT. Regarding the mixed state, we must be careful about what we mean by gaped and gapless for the mixed state. The gap that we are considering here is the energy of the least excited mode on the double state. The terminology “gapless” means that there is an excitation with vanishing energy for $L\to\infty$ in the bulk. In this sense, this gap is different from the one recently proposed in [23, 28, 53], where the gap is defined as the inverse correlation length of a quantum information theoretic correlation.
@@ IV.3 Physical observables
The first ones are $\sigma$-spin correlators to identify the $Z^{\sigma}_{2}$ SWSSB state. The one of them is the partial sum of Rényi-2 correlator,
@@
Another observable is the partial sum of the $\sigma^{z}\sigma^{z}$-correlator of the upper chain,
@@ IV.4 Numerical results
V Detailed comments on co-existance of ASPT and SWSSB
In the previous section, we showed that the double ASPT state emerges as the strength of decoherence increases beyond the phase transition point. In that phase, ASPT and SWSSB coexist as observed by the Rényi-2 correlators, Rényi-2 string order and EE. Here, we discuss the relationship between this observation and previous works that studied the ASPT and SWSSB [23, 28] as the above result may seem odd. There are two concerns.
The first one concerns classification of ASPT. In Ref. [23], from the point of view of the pure state SPT state, the ASPT state is classified by symmetrically invertible property. For the present model, we expect that states under weak decoherence inherit this property as the initial state is the genuine SPT cluster state. As the result of Fig. 2 shows, the Rényi-2 correlator behaves as $C^{\rm II}_{\sigma^{z}\sigma^{z}}(i,j)\to 0$ for the infrared (IR) limit $L,|i-j|\to\infty$. Provided that the fidelity correlator [28] and Rényi-2 correlator exhibit qualitatively the same behavior in the IR limit, SWSSB does not occur there as discussed in  [28], and this weakly decohered phase is an ASPT state of the above classification [23]. However, as the strength of decoherence increases, a phase transition occurs and the double ASPT state appears. As a result of this transition, the resultant state may lose the symmetrically invertible property. In fact, the numerical result $C^{\rm II}_{\sigma^{z}\sigma^{z}}(i,j)\to c\neq 0$ in Fig. 2 indicates the appearance of SWSSB. As Ref. [28] showed, mixed states with SWSSB is not symmetrically invertible. From this observation, we think that the double ASPT state exhibits SWSSB and is not symmetrically invertible, indicating that the double SPT state is not to be classified by the scheme of Ref. [23] (a similar observation appeared in [40]). None the less, we call the strongly-decohered state double ASPT state, as it exhibits the Rényi-2 string order as well as EE, which indicates non-trivial edge modes. Detailed investigation and deep understanding of this interesting phenomenon are future problem.
Second one concerns “gap” of the system. For a density matrix, it is generally difficult to define the gap. However as showed by Ref. [23], the ASPT can have a gap in the sense of the inverse of the correlation length of a quantum information theoretic correlation [53], while a SWSSB has no gap in this sense shown in Ref. [28]. This discrepancy should be clarified. Here, we should comment again on the meaning of gap in this work, that is, we focus on the doubled purified (un-normalized) state defined on the doubled Hilbert space. Gaped state in this formalism is the one that any excitation on the doubled decohered state (including the multiplet coming from the SWSSB) has a finite energy gap for $L\to\infty$. Thus, this notion of gap does not necessarily equate to those in [28, 53]. One interesting question is whether the double ASPT numerically observed here is an exceptional example as the ASPT proposed in [22, 23] or not. This theoretical question can be future work.
VI Summary and conclusion
@@
In addition, we focus on the partial sum of the Rényi-2 correlator, $\chi^{\rm II}$. This quantity can be related to the the fidelity susceptibility [58]. Recently, the fidelity susceptibility is further related to the fluctuation-dissipation theorem [58]. It is an interesting problem to investigate the relationship between $\chi^{\rm II}$ and the fluctuation-dissipation theorem in the mixed phases that we found in this work.